OpenStudy (anonymous):

If cos(alpha) = 2/3 then what is sin(alpha) and tan(alpha)

OpenStudy (slaaibak):

Exactly the same as the other question. Use the following identities: \[\cos^2 x + \sin^2 x = 1\] and \[\tan x = {\sin x \over \cos x}\]

OpenStudy (mertsj):